heathergreen:

What is the value of x in the equation Three-fourths (one-fourth x + 8) minus (one-half x + 2) = StartFraction 3 Over 8 EndFraction (4 minus x) minus one-fourth?

jhonyy9:

\[\frac{ 3 }{ 4 }\left( \frac{ 1 }{ 4 }x +8\right) - \left( \frac{ 1 }{ 2 }x +2 \right) = \frac{ 3 }{ 8 }(4 - x) - \frac{ 1 }{ 4 }\] so like a first step distribute the parentheses